# Re-visiting Skip-Gram Negative Sampling: Dimension Regularization for More Efficient Dissimilarity Preservation in Graph Embeddings

### v1

@@ 2.1. Dimension Regularization for Spectral Embeddings
In the case of Adjacency Spectral Embeddings (ASE) (Daniel L. Sussman and Priebe, 2012), which are equivalent to taking the leading eigenvectors of the adjacency matrix, the matrix $S$ is the adjacency matrix $A\in\{0,1\}^{n\times n}$. For convenience, we define $P$ and $N$ for an individual node $i$, where the full function simply sums over all nodes (e.g., $L\left(X,S\right)=\sum_{i\in V}L\left(X,S,i\right)$):
(2) $L_{ASE}(X,S,i)$ $=\|S_{i}-XX_{i}\|_{2}^{2},$
(3) $P_{ASE}(X,S,i)$ $=-2\sum_{j\in\{j^{\prime}|S_{ij^{\prime}}=1\}}X_{i}^{T}X_{j}+\|S\|_{F}^{2},$
(4) $N_{ASE}(X,S,i)$ $=\|XX_{i}\|_{2}^{2}.$
@@
(5) $X_{i}^{t+1}=X_{i}^{t}-2\eta\sum_{i^{\prime}\in V}\left(X_{i}^{T}X_{i^{\prime}}\right)X_{i^{\prime}}~{}~{}\forall i\in V.$
@@
With Proposition 2.1, we can re-interpret the gradient descent updates in Equation (5) as collectively repelling dimensions. The gradient update can now be written in terms of dimensions:
(6) $X_{.j}^{t+1}=X_{.j}^{t}-2\eta\sum_{j^{\prime}\in[d]}\left(X_{.j}^{T}X_{.j^{\prime}}\right)X_{.j^{\prime}}~{}~{}\forall j\in[d].$
@@ 2.2. Dimension Regularization for Skip-Gram Embeddings
We now introduce a dimension-based approach for skip-gram embeddings. For skip-gram embeddings, the similarity matrix is defined such that $S_{ij}=1$ if node $j$ is in the neighborhood of $i$. For first-order LINE, the neighborhood for node $i$ is simply all nodes connected to $i$ whereas for node2vec, the neighborhood is defined as all nodes reached via a biased random walk initiated at $i$. The skip-gram (SG) loss functions can be decomposed as:
(7) $L_{SG}(X,S,i)$ $=-\sum_{j\in V}S_{ij}\log\sigma\left(X_{i}^{T}X_{j}\right)+(1-S_{ij})\log\sigma\left(-X_{i}^{T}X_{j}\right),$
(8) $P_{SG}(X,S,i)$ $=-\sum_{j\in\{j^{\prime}|S_{ij^{\prime}}=1\}}\log\sigma\left(X_{i}^{T}X_{j}\right),$
(9) $N_{SG}(X,S,i)$ $=-\sum_{j\in\{j^{\prime}|S_{ij^{\prime}}=0\}}\log\sigma\left(-X_{i}^{T}X_{j}\right).$
@@
The constriction $\mathcal{C}\in(0,1)$ of an embedding matrix $X$ is defined as: $\mathcal{C}=\min_{i,j\in n\times n}\sigma\left(X_{i}^{T}X_{j}\right)$.
Geometrically, the embedding constriction is maximized when embeddings are radially squeezed and growing in magnitude, that is, collapsed. Proposition 2.3 states that if we remove $N_{SG}$ altogether and the embeddings are initialized with sufficiently small norm and learning rate, the degenerate collapse will inevitably arise during the course of gradient descent. In the context of graph neural networks, Proposition 2.3 provides conditions under which embedding oversmoothing is guaranteed. In Appendix A.2.1, we validate Proposition 2.3 by showing that embeddings inevitably collapse if only attraction updates are applied for various synthetic sparse graphs.
@@
As the Euclidean norm of the initial embeddings and the learning rate approach zero, then for any constriction threshold $c\in(0,\infty)$ there exists a time step $t$ such that after applying gradient descent on $P_{SG}$ for $t$ epochs the constriction $\mathcal{C}\geq c$.
The proof sketch for Proposition 2.3 is as follows: as the embeddings are initialized closer to the origin, gradient descent on $P_{SG}$ approaches gradient descent on the matrix completion loss function: $\|\mathbf{1}_{S>0}\odot\left(\vec{1}\vec{1}^{T}-\frac{1}{2}XX^{T}\right)\|_{F}^{2}$, where $\mathbf{1}$ is the indicator matrix for whether entries of $S$ are positive. From Gunasekar et al. (2017), gradient descent implicitly regularizes matrix completion to converge to the minimum nuclear solution; this implicit regularization drives all dot products to be positive, not just pairs of embeddings corresponding to connected nodes.
@@
If all embeddings have norms that are neither infinitely large or vanishingly small and the embedding constriction $\mathcal{C}>1/2$, then, as the number of nodes in a sparse graph grows to infinity, the gradients of $\nabla N_{SG}$ and $\nabla N^{\prime}_{SG}$ converge:
@@
For a single node $i$, performing gradient descent on Equation (10) results in the following update:
(12) $X_{i}^{t+1}=\left(1-\eta\sigma\left(\|X_{i}^{t}\|^{2}\right)\right)X_{i}^{t}-\eta\sum_{i^{\prime}\in V}\sigma\left(\left(X_{i}^{t}\right)^{T}X_{i^{\prime}}^{t}\right)X_{i^{\prime}}^{t}.$
The right-hand term in the gradient update repels node $i$ from all other nodes where the repulsion is proportional to the similarity between the node embeddings. In Proposition 2.5, we show a connection between minimizing $N^{\prime}_{SG}$ and centering the dimensions at the origin. For intuition, observe that if all pairs of nodes are highly similar, the gradient update in Equation (12) is approximately equal to subtracting the column means scaled by a constant ($2\eta\left(X^{T}\vec{1}\right)$). This is equivalent to performing gradient descent on a dimension regularizer that penalizes non-zero dimension means,
@@
Let $R$ be the dimension regularizer defined in Equation (13) that penalizes embeddings centered away from the origin and $n\gg d$. Then, as the constriction increases beyond zero, the difference between performing gradient descent on $R$ versus $N^{\prime}_{SG}$ vanishes.
We note that our result establishing a connection between the skip-gram loss and origin-centered dimensions is analogous to the finding in Wang and Isola (2022) connecting the InfoNCE loss with embeddings uniformly distributed on unit hyperspheres.
@@ 2.2.1. Comparison with Skip-Gram Negative Sampling
Skip-gram negative sampling (SGNS) offers an efficient stochastic approximation to the gradient update in Equation (12). Furthermore, the SGNS proceedure provides a tunable way to bias the gradients—via non-uniform sampling—in a manner that has been seen to empirically improve the utility of the resulting embedding in downstream tasks (Mikolov et al., 2013). Instead of repelling node $i$ from all other $n-1$ nodes, SGNS repels $i$ from a sample of $k$ nodes where the nodes are sampled according to a distribution $P_{\alpha}$ over all nodes, optimizing the following objective:
@@
In aggregate, SGNS reduces the gradient time complexity from $\mathcal{O}\left(n^{2}\right)$ to $\mathcal{O}\left(kn\right)$. With Proposition 2.5, instead of reducing the number of pairwise embedding repulsions, we focus on dimension operations, reducing the time complexity from linear in $n$ to linear in $d$.
As mentioned, SGNS embeddings can be tuned by the choice of the non-uniform sampling distribution, where in graph embedding contexts the distribution $P_{\alpha}$ is typically sampling nodes proportional to their degreeα, with $\alpha=3/4$. An optimization-based intuition for this choice is that a degree-based non-uniform distribution prioritizes learning the embeddings of high-degree nodes, but we emphasize that the specific choice of $\alpha=3/4$ is typically motivated directly based on improved empirical performance in downstream tasks.
While beyond the scope of the present work, we briefly note that our dimension regularization framework is immediately amenable to introducing an exactly analogous tuning opportunity. We can simply replace the regularization in Equation (15) with
@@
where $\vec{p}$ is a normalized weight vector that biases the negative update in exact correspondence to the probabilities of each node in $P_{\alpha}$. In our later simulations, we focus our efforts on the uniform case of $\vec{p}=\vec{1}$, i.e., the regularizer in Equation (15).
@@ 3. Algorithm Augmentation to Replace SGNS
We now propose a generic algorithm augmentation framework that directly replaces SGNS with dimension regularization. We instantiate this algorithm augmentation for LINE and node2vec but note that the framework is applicable to any graph embedding algorithm using SGNS.
Our augmentation modifies existing algorithms using SGNS in two ways. First, the augmentation prioritizes the positive function $P$, that is preserving similarity when possible. Not only does prioritizing observed edges increase efficiency, but it is also desirable given the fact that real-world graph data frequently have missing edges (Newman, 2018; Young et al., 2021). We prioritize the observed edges by defaulting to performing gradient descent on $P$.
The second augmentation replaces SGNS with gradient descent on $R$, a dimension regularizer that induces mean-centered dimensions, as introduced in Equation (13). The regularizer directly replaces repelling nodes embedding from negative samples, and, if the embeddings exhibit high constriction after repeated gradient updates to $P$, $\nabla R$ approximates $\nabla N^{\prime}_{SG}$ as established in Proposition 2.5.
@@
(16) $X^{t+1}=\begin{cases}X^{t}-\eta\nabla P_{SG}\left(X^{t}\right)&t\%n_{\text{negative}}\neq 0,\\ X^{t}-\eta\left[\nabla P_{SG}\left(X^{t}\right)+\lambda\nabla R\left(X^{t}\right)\right]&t\%n_{\text{negative}}=0,\end{cases}$
@@
Figure 2 visualizes our algorithm augmentation with a toy stochastic block model, where we compare SGNS and dimension regularization side-by-side. In both cases, we begin by randomly initializing embeddings around the origin and then repeatedly (over 100 epochs) attracting the embeddings of similar nodes. The attraction drives the embeddings away from the origin and causes the two blocks to remain entangled. Then, we apply a repulsion force, using SGNS in the top row and dimension mean regularization in the bottom row, which pulls the embeddings back to the origin. In the case of dimension regularization, once we resume applying positive updates, the embeddings for the two blocks begin to separate, indicating effective repulsion.
Figure 2. We use a toy two-block SBM example to summarize how we bypass skip-gram negative sampling with dimension regularization. We introduce an algorithm augmentation that prioritizes attracting embeddings of similar nodes together (Step 2), where similarity is the sigmoid of embedding dot products. Eventually, to prevent all pairs of embeddings from becoming similar, the dimension regularization re-centers the embeddings around the origin, increasing node contrast (Step 3). Attraction updates are then again repeatedly applied, but the blocks are more distinguishable post regularization (Step 4).
@@ 3.1. Instantiation for LINE and node2vec
In Algorithm 1 we include the pseudo-code for the augmented versions of augmented LINE and node2vec, utilizing the framework in Equation (16). The commented out pseudo-code corresponds to the old negative update code that is removed and replaced by the new negative update code. The below implementation also optimizes the embeddings through batches of positive edges.
@@
Input: $G,n,d,p,q,\text{num\_batches},\text{batch\_size},\lambda,\eta,n_{\text{negative}}$
@@
[…]
         $i,j\leftarrow\texttt{sample\_uniform\_pair}(walks)$
         $X^{t+1}_{i}\leftarrow X^{t}_{i}+\eta\sigma\left(-\langle X^{t}_{i},X^{t}_{j}\rangle\right)X^{t}_{j}$triangleright positive update
         $X^{t+1}_{j}\leftarrow X^{t}_{j}+\eta\sigma\left(-\langle X^{t}_{i},X^{t}_{j}\rangle\right)X^{t}_{i}$
@@
              $X^{t+1}_{i}\leftarrow X^{t}_{i}-\eta\sigma\left(-\langle X^{t}_{i},X^{t}_{j^{\prime}}\rangle\right)X^{t}_{j^{\prime}}$
              $X^{t+1}_{j^{\prime}}\leftarrow X^{t}_{j^{\prime}}-\eta\sigma\left(-\langle X^{t}_{i},X^{t}_{j^{\prime}}\rangle\right)X^{t}_{i}$
@@
         if $\text{idx}~{}~{}\%~{}~{}n_{\text{negative}}==0$ then $\triangleright$ new negative update
              for $j\in\{1,\dots,d\}$ do
                  $X^{t+1}_{.j}\leftarrow X^{t}_{.j}-\frac{\lambda}{n}\sum_{i}X_{ij}\vec{1}$

### v2

@@ 3.1. Dimension Regularization for Spectral Embeddings
In the case of Adjacency Spectral Embeddings (ASE) (Daniel L. Sussman and Priebe, 2012), which are equivalent to taking the leading eigenvectors of the adjacency matrix, the matrix $S$ is the adjacency matrix $A\in\{0,1\}^{n\times n}$.
(2) $L_{ASE}(X,S)$ $=\|S-XX^{T}\|_{2}^{2},$
(3) $P_{ASE}(X,S)$ $=-2\sum_{i,j\mid S_{ij}=1}X_{i}^{T}X_{j}+\|S\|_{F}^{2},$
(4) $N_{ASE}(X,S)$ $=\|XX^{T}\|_{F}^{2}.$
@@
(5) $X_{i}^{t+1}=X_{i}^{t}-\eta\sum_{i^{\prime}\in V}\left(X_{i}^{T}X_{i^{\prime}}\right)X_{i^{\prime}}~{}~{}\forall i\in V.$
@@
With Proposition 3.1, we can re-interpret the gradient descent updates in Eq. (5) as collectively repelling dimensions. The gradient update can now be written in terms of dimensions:
(6) $X_{.j}^{t+1}=X_{.j}^{t}-\eta\sum_{j^{\prime}\in[d]}\left(X_{.j}^{T}X_{.j^{\prime}}\right)X_{.j^{\prime}}~{}~{}\forall j\in[d].$
@@ 3.2. Dimension Regularization for Skip-Gram Embeddings
We now introduce a novel dimension-based approach for skip-gram embeddings. For skip-gram embeddings, the similarity matrix is defined such that $S_{ij}=1$ if node $j$ is in the neighborhood of $i$. For first-order LINE, the neighborhood for node $i$ is simply all nodes connected to $i$, while for node2vec, the neighborhood is defined as all nodes within the context of $i$ on a random walk. The skip-gram (SG) loss function can be decomposed as:
(7) $L_{SG}(X,S)$ $=-\sum_{i,j}S_{ij}\log\sigma\left(X_{i}^{T}X_{j}\right)+(1-S_{ij})\log\sigma\left(-X_{i}^{T}X_{j}\right),$
(8) $P_{SG}(X,S)$ $=-\sum_{i,j\mid S_{ij}=1}\log\sigma\left(X_{i}^{T}X_{j}\right),$
(9) $N_{SG}(X,S)$ $=-\sum_{i,j\mid S_{ij}=0}\log\sigma\left(-X_{i}^{T}X_{j}\right).$
@@
The constriction $\mathcal{C}$ of an embedding matrix $X$ is defined as: $\mathcal{C}=\min_{i,j\in[n]\times[n]}X_{i}^{T}X_{j}$.
Geometrically, the embedding constriction is maximized when embeddings are radially squeezed and growing in magnitude, that is, collapsed. Proposition 3.3 states that if we remove $N_{SG}$ altogether and the embeddings are initialized with sufficiently small norm and learning rate, the degenerate collapse will inevitably arise during gradient descent.
@@
As the Euclidean norm of the initial embeddings and the learning rate approach zero, then in the process of performing gradient descent on $P_{SG}$, there exists a step $t$ such that for all gradient updates after $t$, the constriction $\mathcal{C}$ is positive and monotonically increasing.
The proof sketch for Proposition 3.3 is as follows: as the embeddings are initialized closer to the origin, gradient descent on $P_{SG}$ approaches gradient descent on the matrix completion loss function: $\|\mathbf{1}_{S>0}\odot\left(\mathbf{1}-\frac{1}{2}XX^{T}\right)\|_{F}^{2}$, where $\mathbf{1}_{S>0}$ is the indicator matrix for whether entries of $S$ are positive. From Gunasekar et al. (2017), gradient descent implicitly regularizes matrix completion to converge to the minimum nuclear solution; this implicit regularization drives all dot products to be positive, not just pairs of embeddings corresponding to connected nodes.
Approaching dimension regularization.
@@
If all embeddings have norms that are neither infinitely large or vanishingly small and the embedding constriction $\mathcal{C}>0$, then, as the number of nodes in a sparse graph grows to infinity, the gradients of $\nabla N_{SG}$ and $\nabla N^{\prime}_{SG}$ converge:
@@
For a single node $i$, performing gradient descent on Eq. (10) results in the following update:
(12) $X_{i}^{t+1}=X_{i}^{t}-\eta\sum_{i^{\prime}\in V}\sigma\left(\left(X_{i}^{t}\right)^{T}X_{i^{\prime}}^{t}\right)X_{i^{\prime}}^{t}.$
The right-hand term in the gradient update repels node $i$ from all other nodes where the repulsion is proportional to the similarity between the node embeddings. In Proposition 3.5, we show a connection between minimizing $N^{\prime}_{SG}$ and centering the dimensions at the origin. For intuition, observe that if all pairs of nodes are highly similar, the gradient update in Eq. (12) is approximately equal to subtracting the column means scaled by a constant ($2\eta\left(X^{T}\vec{1}\right)$). This is equivalent to performing gradient descent on a dimension regularizer that penalizes non-zero dimension means,
@@
Let $R$ be the dimension regularizer defined in Eq. (13) that penalizes embeddings centered away from the origin and $n\gg d$. Then, as the constriction increases beyond zero, the difference between performing gradient descent on $R$ versus $N^{\prime}_{SG}$ vanishes.
We note that our result establishing a connection between the skip-gram loss and origin-centered dimensions is related to the finding in Wang and Isola (2020) connecting the InfoNCE loss with embeddings uniformly distributed on unit hyperspheres; however, as the SG and InfoNCE losses are distinct, our theoretical contribution in Proposition 3.5 is novel.
@@ 3.2.1. Comparison with Skip-Gram Negative Sampling
Skip-gram negative sampling (SGNS) offers an efficient stochastic approximation to the gradient update in Eq. (12). Furthermore, the SGNS procedure provides a tunable way to bias the gradients—via non-uniform sampling—in a manner that has been seen to empirically improve the utility of the resulting embedding in downstream tasks (Mikolov et al., 2013). Instead of repelling node $i$ from all other $n-1$ nodes, SGNS repels $i$ from a sample of $k$ nodes where the nodes are sampled according to a distribution $P_{\alpha}$ over all nodes, optimizing the following objective:
@@
In aggregate, SGNS reduces the repulsion time complexity from $\mathcal{O}\left(n^{2}\right)$ to $\mathcal{O}\left(kbn\right)$ vector additions per epoch. $b$ is the average number of positive pairs per node; for LINE, $b=m/n$, and for node2vec, $b$ is the product of the context size and random-walk length. With Proposition 3.5, we reduce the time complexity to $\mathcal{O}(d)$ vector additions per epoch.
As mentioned, SGNS embeddings can be tuned by the choice of the non-uniform sampling distribution, where in graph embedding contexts the distribution $P_{\alpha}$ is typically sampling nodes proportional to their degreeα, with $\alpha=3/4$. An optimization-based intuition for this choice is that a degree-based non-uniform distribution prioritizes learning the embeddings of high-degree nodes, but we emphasize that the specific choice of $\alpha=3/4$ is typically motivated directly based on improved empirical performance in downstream tasks (Mikolov et al., 2013).
We briefly note that our dimension-mean regularizer is immediately amenable to introducing an analogous tuning opportunity. We can simply replace the regularization in Eq. (13) with
@@
where $\vec{p}$ is a normalized weight vector that biases the negative update in exact correspondence to the probabilities of each node in $P_{\alpha}$. In our subsequent simulations, we focus our efforts on the uniform case of $\vec{p}=\vec{1}$, i.e., the regularizer in Eq. (13).
@@ 4. Algorithm Augmentation to Replace SGNS
We now propose a generic algorithm augmentation framework that directly replaces SGNS with dimension regularization. We instantiate this algorithm augmentation for LINE and node2vec, but note that the framework is applicable to any graph embedding algorithm using SGNS.
Our augmentation modifies existing algorithms using SGNS in two ways. First, the augmentation prioritizes the positive function $P$, that is, preserving similarity when possible. Not only does prioritizing observed edges increase efficiency, but it is also desirable given the fact that real-world graph data frequently have missing edges. We prioritize the observed edges by defaulting to performing gradient descent on $P$.
Second, our augmentation achieves repulsion via gradient descent on $R$, the dimension-mean regularizer introduced in Eq. (13). As constriction increases from optimizing only $P$, dimension-mean regularization approximates $\nabla N^{\prime}_{SG}$ per Proposition 3.5.
@@
(16) $X^{t+1}=\begin{cases}X^{t}-\eta\nabla P_{SG}\left(X^{t}\right)&t~{}\%~{}n_{\text{negative}}\neq 0,\\ X^{t}-\eta\left[\nabla P_{SG}\left(X^{t}\right)+\lambda\nabla R\left(X^{t}\right)\right]&t~{}\%~{}n_{\text{negative}}=0,\end{cases}$
@@
Figure 2 visualizes our algorithm augmentation with a Stochastic Block Model (SBM). We begin by randomly initializing embeddings around the origin and then repeatedly (over 100 epochs) attract the embeddings of similar nodes. The attraction drives the embeddings away from the origin and causes the two blocks to remain entangled. Then, we repel dissimilar pairs via dimension-mean regularization, which pulls the embeddings back to the origin. Once we resume applying positive updates, the embeddings for the two blocks begin to separate, indicating effective repulsion.
[…]
[Toy example of our algorithm augmentation]On the left is a toy two-block SBM graph. The right side of the figure shows the embedding of the SBM graph at four snapshots. The snapshots show: 1) random initialization at the origin 2) approaching collapse away from the origin after positive updates 3) retraction to the origin from dimension-mean regularization 4) expansion from the origin but with the two blocks separated in embedding space.
Figure 2. We use a two-block Stochastic Block Model (SBM) example to summarize how we bypass skip-gram negative sampling with dimension regularization. We introduce an algorithm augmentation that prioritizes attracting embeddings of similar nodes together (Step 2), where similarity is the sigmoid of embedding dot products. Eventually, to prevent all pairs of embeddings from becoming similar, the dimension regularization re-centers the embeddings around the origin, increasing node contrast (Step 3). Attraction updates are then again repeatedly applied, but the blocks are more distinguishable following dimension regularization (Step 4).
@@ 4.1. Instantiation for LINE and node2vec
In Algorithm 1 we include the pseudo-code for the augmented versions of LINE and node2vec, utilizing the framework in Eq. (16). The loop labeled “old negative update” shows where SGNS would have occurred. When instantiated for LINE and node2vec, we implement our algorithm augmentation with batched gradient descent, where batches are sets of positive pairs. Dimension regularization is applied every $n_{\text{negative}}$ batches, and the regularization is applied to the entire embedding matrix $X$.
@@
Input: $G,n,d,p,q,\text{num\_epochs},\text{batch\_size},\lambda,\eta,n_{\text{negative}}$
$t\leftarrow 0$
@@
         for $i,j\in$ batch do
              $X^{t+1}_{i}\mathrel{+}=~{}\eta\sigma\left(-\langle X^{t}_{i},X^{t}_{j}\rangle\right)X^{t}_{j}$triangleright positive update
              $X^{t+1}_{j}\mathrel{+}=~{}\eta\sigma\left(-\langle X^{t}_{i},X^{t}_{j}\rangle\right)X^{t}_{i}$
@@
                  $X^{t+1}_{i}\mathrel{-}=~{}\eta\sigma\left(-\langle X^{t}_{i},X^{t}_{j^{\prime}}\rangle\right)X^{t}_{j^{\prime}}$
                  $X^{t+1}_{j^{\prime}}\mathrel{-}=~{}\eta\sigma\left(-\langle X^{t}_{i},X^{t}_{j^{\prime}}\rangle\right)X^{t}_{i}$
@@
         $t\leftarrow t+1$
@@
     if $t~{}~{}\%~{}~{}n_{\text{negative}}==0$ then $\triangleright$ new negative update
         $X^{t+1}\mathrel{-}=~{}\frac{\lambda}{n}\mathbf{1}X^{t}$